The average (arithmetic mean) of a set of 5 numbers is 9.6. When a sixth number is added to the set, the average of the set becomes 9. What is the sixth number?
Answer by alicealc(293)

(5 * 9.6 + x)/6 = 9
(48 + x) = 6 * 9
48 + x = 54
x = 54 - 48 = 6

so, the sixth number is 6